Using the the concept of rank of the matrix find whether the following system of equations

are consistent or inconsistent,

i.

4y + z = 0

12x - 5y - 3z = 34

-6x + 4z = 8


1
Expert's answer
2021-06-09T08:44:28-0400

i. "Ax=b" is inconsistent (i.e., no solution exists) if and only if "\text{rank}[A]<\text{rank}[A|b]."


ii. "Ax=b" has an unique solution if and only if "\text{rank}[A]=\text{rank}[A|b]=n."


iii. "Ax=b" has infinitely many solutions if and only if "\text{rank}[A]=\text{rank}[A|b]<n."



"A=\begin{bmatrix}
 0 & 4 & 1 \\
 12 & -5 & -3 \\
 -6 & 0 & 4
\end{bmatrix}"

Swap rows 1and 2


"\begin{bmatrix}
 12 & -5 & -3 \\
 0 & 4 & 1 \\
 -6 & 0 & 4
\end{bmatrix}"

"R_3=R_3+(1\/2)R_1"

"\begin{bmatrix}
 12 & -5 & -3 \\
 0 & 4 & 1 \\
 0 & -5/2 & 5/2
\end{bmatrix}"

"R_3=R_3+(5\/8)R_2"


The rank of a matrix is the number of nonzero rows in the reduced matrix, so the rank is 3.


Since "\text{rank}[A]=\text{rank}[A|b]=3=n," then "Ax=b" has an unique solution.

Therefore, the given system of equations is consistent.